# Large Language Models for cross-language code clone detection

### v1

@@ Abstract.
With the involvement of multiple programming languages in modern software development, cross-lingual code clone detection has gained traction with the software engineering community. Numerous studies have explored this topic, proposing various promising approaches. Inspired by the significant advances in machine learning in recent years, particularly Large Language Models (LLMs), which have demonstrated their ability to tackle various tasks, this paper revisits cross-lingual code clone detection.
We investigate the capabilities of four (04) LLMs and eight (08) prompts for the identification of cross-lingual code clones. Additionally, we evaluate a pre-trained embedding model to assess the effectiveness of the generated representations for classifying clone and non-clone pairs. Both studies (based on LLMs and Embedding models) are evaluated using two widely used cross-lingual datasets, XLCoST and CodeNet. Our results show that LLMs can achieve high F1 scores, up to 0.98, for straightforward programming examples (e.g., from XLCoST). However, they not only perform less well on programs associated with complex programming challenges but also do not necessarily understand the meaning of code clones in a cross-lingual setting. We show that embedding models used to represent code fragments from different programming languages in the same representation space enable the training of a basic classifier that outperforms all LLMs by $\sim$2 and $\sim$24 percentage points on the XLCoST and CodeNet datasets, respectively. This finding suggests that, despite the apparent capabilities of LLMs, embeddings provided by embedding models offer suitable representations to achieve state-of-the-art performance in cross-lingual code clone detection.
Large Language Model, Embedding Model, Prompt Engineering, Cross-Language Pairs, Code Clone Detection
@@ 1. Introduction
Copy-pasting a code fragment with or without change is common practice in software development, and this action produces code clones (Roy and Cordy, [n. d.]). It reduces the quality of the code, making maintenance costly, and can be a source of bugs (Juergens et al., 2009; Roy and Cordy, [n. d.]). Research shows that we can find between 5% to 23% of clones in a software system (White et al., 2016; Koschke, [n. d.]). We distinguish four types of clones ranging from type 1 to type 4 (Keller et al., 2022). Types 1, 2, and 3 clones are generally classified as syntactic clones, and type 4 is referred to as semantic clones. Syntactic clones occur when there are textual similarities, while semantic clones represent functional similarity. Type 4 is the most complicated clone and the hardest to detect (Dou et al., 2023).
@@
We often retrieve code clones within the same programming language in software development. However, in modern and complex software development, the multiplicity of software components implies using several programming languages to solve the challenges (Kochhar et al., 2016). For instance, developers can use C, C++, Java, or Kotlin for Android applications and Objective-C for iPhone applications. Intentionally, developers may introduce clones for building identical systems for different platforms. Also, the development of those systems is, for the most part, collaborative. So, if an expert in a certain type of programming language changes one part of the software for a given language, the exact functional change must be done on the other variants. This is more resource and time-consuming than modifying a software system based on a single language, as it requires prior knowledge of the system architecture and an understanding of the code changes performed by the first developer (Nafi et al., 2019). To manage cross-lingual systems in an easy, time-effective, and cost-effective way, developers need an automatic system that can detect clones of various languages at the same time(Nafi et al., 2019).
@@
The literature includes several approaches and tools for code clone detection. For instance, CCFinder  (Kamiya et al., 2002), Deckard  (Jiang et al., 2007), NiCad  (Roy and Cordy, 2008), SourcererCC (Sajnani et al., 2016), NIL (Nakagawa et al., 2021), and OREO (Saini et al., 2018) have been developed for type 1, 2, and 3 clones detection within the same programming language. For semantic clones within the same language, RtvNN  (White et al., 2016), CDLH  (Wei and Li, 2017), DeepSim (Zhao and Huang, 2018), and FA-AST  (Wang et al., 2020) have been proposed. Concerning cross-lingual clones detection, tools such as LICCA (Vislavski et al., 2018), CLCMiner (Cheng et al., 2017), CLCDSA (Nafi et al., 2019), C4 (Tao et al., 2022), TCCCD (Fang et al., 2023) , CCT-Code (Sorokin et al., 2023), AdaCCD  (Du et al., 2024) and  (Khajezade et al., 2024) have been introduced. They rely on deep learning techniques to capture the syntactic and semantic relationships between different parts of the source code. CLCDSA employed on-the-fly and deep neural network-based approaches using Abstract Syntax Trees (ASTs), a fundamental data structure in computer science for representing code structure, to extract features for cross-lingual code clone detection. On the fly approach measures cosine similarity between the two matrices generated by extracting values of the features from source code fragments while the deep neural network-based approach learns automatically the values of features from data and detects cross-lingual clones. TCCCD employed the UniXcoder  (Guo et al., 2022) pre-trained model to map code clones from different languages into vector space and fine-tuned the model using triplet learning. Sorokin et al. proposed an approach using cross-consistency training (CCT) for language models training on source code in different programming languages and presented the CCT-LM model, initialized with GraphCodeBERT and fine-tuned with CCT to retrieve semantically similar codes given a code as the query. AdaCCD leverages language-agnostic representations from pre-training programming language models (GraphCodeBERT and CodeBERT) and contrastive learning for cross-lingual adaptation to detect code clones.
@@
In this paper, we investigate Large Language Models (LLMs) and Embedding Models’ (EM) capabilities for cross-lingual code clone detection. LLMs performances being influenced by prompt engineering (Lin, 2024; Bai et al., [n. d.]; Yin et al., 2024), we first design various different prompts that we applied on four LLMs (Falcon-7B-Instruct  (Almazrouei et al., 2023), LLAMA2-Chat-7B (Touvron et al., 2023), Starchat-$\beta$  (Tunstall et al., 2023), and GPT-3.5-Turbo). The design prompts include two simple prompts: one that asks the LLMs to directly provide ”yes/no” answers if the two code snippets are clones or not, and an improved version that includes instructions asking the LLMs to take into account the whole structure and logic of the two code snippets and conclude if they perform a similar task. In addition to those simple prompts, we designed one-step and multi-step Chain of Thought (CoT) (Dou et al., 2023) prompts to evaluate their impact on LLMs’ performances. With these prompts, we want the LLMs to follow step-by-step reasoning to perform the task by taking into account the code similarity, the similar lines, the thinking process, and the code description analysis.
Our second study for cross-lingual code clone detection relies on the use of an embedding model. We select the pre-trained model ‘Text-Embedding-Ada-002’ provided by OpenAI. Given two code snippets, we directly provide them as inputs to the embedding model, which outputs two vectors. To detect whether the provided code snippets are clones or not, we can directly perform a similarity measure using the two vectors and cosine similarity (Pedregosa et al., 2011). This requires the definition of a threshold. In our experiments, we investigate several threshold values. With the embedding vectors, we also conducted additional experiments for clone detection by training custom binary classifiers.
@@
Our experimental results show that GPT-3.5-Turbo achieved an overall F1 score of 0.98 for the XLCoST and 0.75 for CodeNet. It outperforms LLAMA2-Chat-7B, Starchat-$\beta$, and Falcon-7B-Instruct that achieve respectively overall F1 scores of 0.81, 0.76, and 0.41 for the XLCoST dataset and 0.65, 0.54, and 0.38 for the CodeNet dataset. Among all the designed prompts, the “improved simple prompt” is the one that enables all LLMs to achieve their best performance in the task of code clone detection. This underlines that LLMs’ understanding of the notion of clones is narrow but that, once provided with definition guidance, they are relatively able to reason across programming languages.
Regarding the embedding model, we first note that the cosine similarity derived from code snippets’ vectors can be directly used to identify code clones. Indeed, with cross-lingual pairs, the embedding model produces close vectors when the code snippets are clones and distant vectors when they are non-clones. This is observed in  70% of the pairs with a cosine similarity threshold of 60%. We further leverage the vector pairs to train several classifiers for cross-lingual code clone detection. Among those classifiers, the Support Vector Machine classifier (SVM) using a polynomial kernel achieves F1-scores of 0.998 and 0.995, respectively, for XLCoST and CodeNet. In sum, all our evaluations reveal that the embedding model outperforms all LLMs, even though the latter yields satisfactory results when combined with CoT-based prompts.
This work makes several key contributions to the field of cross-lingual code clone detection:
@@
  • We discuss the performance of LLMs by comparing them against the classical approach of machine learning classification based on learned representations of code. Experimental results even hint that LLMs may not even “understand” the meaning of ”clones” in the context of clones.
@@ 3.1. Research questions
  • RQ4: Are LLMs outperforming classical classification models for cross-lingual code clone detection? Before the advent of LLMs, state-of-the-art approaches for code clone detection were employing machine learning-based classification, notably using deep representation learning. We thus compare LLMs against baseline classifiers (e.g. with SVM and kNN) based on Embedding models. The objective is to shed lights into the power of representations vs reasoning complexity in LLMs.
@@ 3.2. Models
We evaluated four LLMs and one embedding model for cross-lingual code clone detection. We chose the popular GPT-3.5-Turbo model from OpenAI due to its prevalence in LLM-based software engineering research. To explore open-source options, we included Falcon-7B-Instruct, LLAMA2-Chat-7B, and Starchat-$\beta$, all trained on multiple programming languages. Unlike prior studies focusing solely on LLMs, we incorporated the Text-embedding-Ada-002 model, also from OpenAI’s web API, to investigate the effectiveness of pre-trained embeddings for this task. We provide detailed descriptions of each model in the following section.
@@
  • Text-Embedding-Ada-002 is a text embedding model created by OpenAI. It is designed for tasks such as text search, text similarity evaluation, and code search. The model generates numerical representations (embeddings) of text or code inputs. Embeddings are low-dimensional, dense vector representations that capture the semantic relationships between words or code elements. This allows computers to more efficiently process and understand the meaning of text and code.
@@ 3.3. Datasets
For our experiments, we consider random subsets of 3000 clone pairs from XLCoST and 3000 clone pairs from CodeNet.
Construction of negative sets. To adequately benchmark the different clone detection approaches, we must create negative samples of non-clone pairs. We thus consider creating negative pairs within and accross programming languages. For the XLCost dataset, each problem has a short description and a unique ID. We therefore mixed the clone pairs having different IDs but from the same programming languages, and then used an LLM to confirm the similarity scores of the new pairs’ descriptions. For only a few cases the similarity score is larger than 50%. Nevertheless, we manually confirmed that the descriptions and the code are actually different.
For the CodeNet dataset, we build clone pairs by considering two steps: first, each fragment in a pair must be related to a different problem than its associated fragment in the pair (based on the ID). Second, the fragments in a pair must come from different programming languages. Overall, built the non-clone pairs such as having 300 code fragments for each of the programming languages.

### v2

@@ Abstract.
With the involvement of multiple programming languages in modern software development, cross-lingual code clone detection has gained traction within the software engineering community. Numerous studies have explored this topic, proposing various promising approaches. Inspired by the significant advances in machine learning in recent years, particularly Large Language Models (LLMs), which have demonstrated their ability to tackle various tasks, this paper revisits cross-lingual code clone detection. We evaluate the performance of five (05) LLMs and eight prompts (08) for the identification of cross-lingual code clones. Additionally, we compare these results against two baseline methods. Finally, we evaluate a pre-trained embedding model to assess the effectiveness of the generated representations for classifying clone and non-clone pairs. The studies involving LLMs and Embedding models are evaluated using two widely used cross-lingual datasets, XLCoST and CodeNet.
Our results show that LLMs can achieve high F1 scores, up to 0.99, for straightforward programming examples. However, they not only perform less well on programs associated with complex programming challenges but also do not necessarily understand the meaning of “code clones” in a cross-lingual setting. We show that embedding models used to represent code fragments from different programming languages in the same representation space enable the training of a basic classifier that outperforms all LLMs by $\sim$1 and $\sim$20 percentage points on the XLCoST and CodeNet datasets, respectively. This finding suggests that, despite the apparent capabilities of LLMs, embeddings provided by embedding models offer suitable representations to achieve state-of-the-art performance in cross-lingual code clone detection.
Cross-Language Pairs, Code Clone Detection, Large Language Model, Prompt Engineering, Embedding Model
ccs: Computing methodologies Machine learning approachesccs: Software and its engineering Software maintenance tools
@@ 1. Introduction
Copy-pasting a code fragment with or without change is common practice in software development. It leads to what is commonly referred to as code clones (Roy and Cordy, [n. d.]). While clones provide some benefits (e.g., faster prototyping, reuse of proven code), they are also known to contribute to reducing the quality of the code, making maintenance costly, and being a source of bugs (Juergens et al., 2009; Roy and Cordy, [n. d.]). Research shows that we can find between 5% to 23% of clones in a software system (White et al., 2016; Koschke, [n. d.]). The literature enumerates four types of clones (Keller et al., 2022). Types 1, 2, and 3 clones are generally classified as syntactic clones, and type 4 is referred to as semantic clones. Syntactic clones occur when there are textual similarities, while semantic clones represent functional similarity. Type 4 is the most complicated clone and the hardest to detect (Dou et al., 2023).
@@
Code clone retrieval is generally studied in the context of program fragments written with the same programming language. However, in modern software development, the multiplicity of software components involves various programming languages used, each for each efficiency in addressing specific programming challenges (Kochhar et al., 2016). Developers may for example intentionally introduce clones for building identical systems for different platforms. Since the development of those systems is, for the most part, collaborative, if an expert in a certain type of programming language changes one part of the software for a given language, the exact functional change must be carried over to the other variants. This is more resource and time-consuming than modifying a software system based on a single language, as it requires prior knowledge of the system architecture and an understanding of the code changes performed by the first developer (Nafi et al., 2019). To manage cross-lingual systems in an easy, time-effective, and cost-effective way, developers need an automatic system that can detect clones of various languages at the same time(Nafi et al., 2019).
The literature includes several approaches and tools for code clone detection. For instance, CCFinder  (Kamiya et al., 2002), Deckard  (Jiang et al., 2007), NiCad  (Roy and Cordy, 2008), SourcererCC (Sajnani et al., 2016), NIL (Nakagawa et al., 2021), and OREO (Saini et al., 2018) have been developed for type 1, 2, and 3 clones detection within the same programming language. For semantic clones within the same language, RtvNN  (White et al., 2016), CDLH  (Wei and Li, 2017), DeepSim (Zhao and Huang, 2018), and FA-AST  (Wang et al., 2020) have been proposed.
@@
Concerning cross-lingual clones detection, tools such as LICCA (Vislavski et al., 2018), CLCMiner (Cheng et al., 2017), CLCDSA (Nafi et al., 2019), OneSpace (El Arnaoty and Servant, 2024), C4 (Tao et al., 2022), TCCCD (Fang et al., 2023) , CCT-Code (Sorokin et al., 2023), AdaCCD  (Du et al., 2024) and  (Khajezade et al., 2024) have been introduced. They rely on deep learning techniques to capture the syntactic and semantic relationships between different parts of the source code. CLCDSA employed on-the-fly and deep neural network-based approaches using Abstract Syntax Trees (ASTs), a fundamental data structure in computer science for representing code structure, to extract features for cross-lingual code clone detection. On the fly approach measures cosine similarity between the two matrices generated by extracting values of the features from source code fragments while the deep neural network-based approach learns automatically the values of features from data and detects cross-lingual clones. OneSpace introduces a novel approach to cross-language code clone detection by mapping programming languages into a shared embedding space using both code and API data. It employs a Siamese network to evaluate the similarity between embedded programs and explores an alternative method using Word Mover’s Distance for efficiency. The approach also examines the influence of factors such as code complexity and training data size on its performance. TCCCD employed the UniXcoder  (Guo et al., 2022a) pre-trained model to map code clones from different languages into vector space and fine-tuned the model using triplet learning. Sorokin et al. proposed an approach using cross-consistency training (CCT) for language models training on source code in different programming languages and presented the CCT-LM model, initialized with GraphCodeBERT and fine-tuned with CCT to retrieve semantically similar codes given a code as the query. AdaCCD leverages language-agnostic representations from pre-training programming language models (GraphCodeBERT and CodeBERT) and contrastive learning for cross-lingual adaptation to detect code clones.
@@
In this paper, we investigate Large Language Models (LLMs), and Embedding Models’ (EM) capabilities for cross-lingual code clone detection. LLMs performances being influenced by prompt engineering (Lin, 2024; Bai et al., [n. d.]; Yin et al., 2024), we first design various different prompts that we applied on five LLMs (Falcon-7B-Instruct  (Almazrouei et al., 2023), LLAMA2-Chat-7B (Touvron et al., 2023), Starchat-$\beta$  (Tunstall et al., 2023), StarCoder2-15b-Instruct (Wei et al., 2024) and GPT-3.5-Turbo). The design prompts include two simple prompts: one that asks the LLMs to directly provide ”yes/no” answers if the two code snippets are clones or not, and an improved version that includes instructions asking the LLMs to take into account the whole structure and logic of the two code snippets and conclude if they perform a similar task. In addition to those simple prompts, we designed one-step and multi-step Chain of Thought (CoT) (Dou et al., 2023) prompts to evaluate their impact on LLMs’ performances. With these prompts, we want the LLMs to follow step-by-step reasoning to perform the task by taking into account the code similarity, the similar lines, the thinking process, and the code description analysis.
Our second study for cross-lingual code clone detection relies on the use of embedding models. We selected three pre-trained models: GraphCodeBERT  (Guo et al., 2020) and UniXcoder  (Guo et al., 2022b), two widely used open-source embedding models for both natural language (e.g., descriptions, comments) and programming languages (source code), and the ‘Text-embedding-3-large’ provided by OpenAI.
For GraphCodeBERT and UniXcoder models, which are also used as baselines, similar to prior work (Fang et al., 2023; Khajezade et al., 2024) we directly leverage their ability to be modified for downstream tasks (in our case, cross-lingual code clone detection).
For the ‘Text-embedding-3-large’ model, given two code snippets, we directly provide them as inputs to the embedding model, which outputs two vectors. To detect whether the provided code snippets are clones or not, we can directly perform a similarity measure using the two vectors and cosine similarity (Pedregosa et al., 2011). This requires the definition of a threshold. In our experiments, we investigate several threshold values. With the embedding vectors, we also conducted additional experiments for clone detection by training custom binary classifiers.
@@
Our experimental results show that GPT-3.5-Turbo achieved an overall F1 score of 0.99 for the XLCoST and 0.82 for CodeNet. It outperforms LLAMA2-Chat-7B, Starchat-$\beta$, StarCoder2-Instruct (Wei et al., 2024) and Falcon-7B-Instruct that achieve respectively overall F1 scores of 0.82, 0.615, 0.595, and 0.48 for the XLCoST dataset and 0.58, 0.52, 0.47, and 0.43 for the CodeNet dataset. Among all the designed prompts, the “improved simple prompt” is the one that enables all LLMs to achieve their best performance in the task of code clone detection. This underlines that LLMs’ understanding of the notion of clones is narrow but that, once provided with definition guidance, they are relatively able to reason across programming languages.
Regarding the embedding model, we first note that the cosine similarity derived from code snippets’ vectors can be directly used to identify code clones. Indeed, with cross-lingual pairs, the embedding model produces close vectors when the code snippets are clones and distant vectors when they are non-clones. This is observed in $\sim$74of the pairs with a cosine similarity threshold of 50%. We further leverage the vector pairs to train several classifiers for cross-lingual code clone detection. Among those classifiers, the Support Vector Machine classifier (SVM) using a polynomial kernel achieves F1-scores of 1 and 0.986, respectively, for XLCoST and CodeNet. In sum, all our evaluations reveal that the embedding model outperforms all LLMs, even though the latter yields satisfactory results when combined with CoT-based prompts.
This research addresses the challenging problem of cross-language clone detection, which requires a deeper semantic understanding of code snippets across diverse syntaxes, structures, and programming styles. By leveraging the capabilities of large language models (LLMs), we explore their potential and limitations in this complex domain and propose strategies, including enhanced prompt designs, to improve their effectiveness. We introduce an approach that combines binary classifiers with code embeddings to effectively capture semantic similarities in cross-language scenarios. This method offers a practical solution for analyzing multilingual codebases, addressing critical challenges in areas such as multilingual code refactoring, security analysis, and software maintenance. By focusing on the application of LLMs to cross-language clone detection, this research contributes to an underexplored area with significant implications for managing large and diverse software ecosystems. This work makes several key contributions to the field of cross-lingual code clone detection:
@@
  • We discuss the performance of LLMs by comparing them against two baselines and the classical approach of machine learning classification based on learned representations of code. Experimental results even hint that LLMs may not even “understand” the meaning of ”clones” in the context of clones.
@@ 3.1. Research questions
  • RQ4: Are LLMs outperforming classification models for cross-lingual code clone detection? Before the advent of LLMs, state-of-the-art approaches for code clone detection were employing machine learning-based classification, notably using deep representation learning. We thus compare LLMs against baseline classifiers (e.g. with SVM and kNN) based on Embedding models. The objective is to shed lights into the power of representations vs reasoning complexity in LLMs.
@@ 3.2. Models
We evaluated five LLMs and one embedding model for cross-lingual code clone detection. The selection process prioritized models fine-tuned for programming languages and code-related tasks, as well as those offering reproducibility and accessibility. From OpenAI, we included the popular GPT-3.5-Turbo model due to its prevalence in software engineering research and its balance of capability and cost-effectiveness, making it suitable for large-scale evaluations. To explore open-source options, we include Falcon-7B-Instruct, Starchat-$\beta$, and LLAMA2-Chat-7B. These models are trained on multiple programming languages and some have been employed in prior studies for code clone detection within the same programming language. Among these, StarCoder2-15b-Instruct was selected for its robust fine-tuning on programming languages, making it particularly well-suited for our task. Additionally, we incorporated the Text-embedding-3-large model, an embedding model from OpenAI’s web API, to investigate the effectiveness of pre-trained embeddings for this task. By including this model, we extended our analysis beyond LLMs to assess the broader potential of embedding-based approaches. we provide detailed descriptions of each model in the following section.
@@
  • StarCoder2-15b-Instruct  (Wei et al., 2024) is the first code LLM trained entirely through a self-aligned process, eliminating the need for human annotations or distilled data from proprietary models. Our open-source pipeline leverages StarCoder2-15B  (Lozhkov et al., 2024) to generate a substantial dataset of instruction-response pairs. These pairs are then used to iteratively refine the model’s capabilities, ensuring a fully transparent and permissive training process.
  • Text-embedding-3-large is a text embedding model created by OpenAI. It is designed for tasks such as search, clustering, recommendations, anomaly detection, diversity measurement, classification text similarity evaluation, and code search. The model generates numerical representations (embeddings) of text or code inputs. Embeddings are low-dimensional, dense vector representations that capture the semantic relationships between words or code elements. This allows computers to more efficiently process and understand the meaning of text and code.
Baselines. We compare the results of the LLMs against two baseline models, chosen for their robust performance in previous research, to strengthen our evaluation.
  • The GraphCodeBERT  (Guo et al., 2021) is a transformer-based model that prioritizes data flow over traditional syntactic structures, such as abstract syntax trees (AST), in its pre-training phase. By leveraging structure-aware approaches, it enhances code representation, making it particularly effective for tasks like code clone detection, translation, refinement, and search. This focus on the intrinsic flow of data within code enables GraphCodeBERT to capture deeper semantic relationships, leading to improved performance across various code understanding tasks.
  • The Unixcoder  (Guo et al., 2022a) is a cross-modal pre-trained model for programming languages that improves upon encoder-decoder frameworks by using mask attention matrices with prefix adapters for fine-grained control. It leverages data such as abstract syntax trees (ASTs) and code comments to enhance source code representation. Pre-trained on nine programming languages, including Java, Ruby, Python, PHP, Javascript, Go, C, C++, and C#, UniXcoder offers enhanced capabilities in representing cloned code, particularly in C++ and C#.
@@ 3.3. Datasets
Construction of sample pairs. To ensure the highest quality samples for our experiments, we implemented a rigorous selection process for both non-clone and clone pairs. Both raw datasets include samples with a problem ID, description, and code snippets in various programming languages. This structure allows us to easily retrieve positive pairs, defined as “two code snippets in different programming languages addressing the same programming problem”. However, it is challenging to build a ground truth set of negative pairs. We postulate that a pair can be negative for two main reasons: (i) the problems addressed by the two code snippets are different; (ii) for the same problem two code snippets actually differ semantically (e.g., one is incorrect). For constructing our dataset, we only select code validated by developers. This means that negative pairs are only those where the problems are different. We therefore propose to use sentence embeddings to compare problem descriptions and build the dataset of negative pairs. Here is the breakdown:
  • For each problem description, we generate the corresponding sentence vector using a state-of-the-art sentence vector generation model (all-mpnet-base-v2  [^5] https://huggingface.co/sentence-transformers/all-mpnet-base-v2). This model has been widely validated for its effectiveness in capturing semantic meaning.
  • To group similar problems, we applied the DBSCAN clustering algorithm to the sentence vectors. Within each cluster, we selected the most complicated problem based on a pre-computed complexity metric. We identify the furthest problem (by vector distance) as a candidate for a negative pair (considering all the problems in the dataset).
  • Negative set selection: We ensure that there are no duplicates or over-representation of specific problems and we prioritize pairing the most complex code.
  • Positive set selection: The remaining problems were used to construct the positive set, with the most complex code selected to represent the clone pairs.
    Before the previously described process, we calculated the cyclomatic complexity of each problem and code sample provided by the two datasets selecting the most complex instances to ensure a representative and challenging dataset.
For our experiments, we used subsets of 3000 clone pairs and 3000 non-clone pairs from each dataset, with Java selected as the primary language cloned alongside four and ten programming languages, respectively, for the XLCoST and, the CodeNet dataset with the same ratio for each programming language.
For the baselines, We employ a similar selection process. However, we restricted our dataset to CodeNet, given its broader coverage of programming languages. From CodeNet, we selected only the programming languages that were compatible with the baseline models. Additionally, we filtered the dataset to include only pairs that met the input size requirements of the baseline models.